How do I express this in exp form?
2Ln(x)+3Ln(y)=2

Answer by stanbon(48535)

express this in exp form?
2Ln(x)+3Ln(y)=2
Ln(x^2) + Ln(y^3) = 2
Ln[x^2*y^3] = 2
x^2y^3 = e^2
That is the the exponential form.
=================================
Cheers,
Stan H.